OpenStudy (anonymous):

A 12-foot ladder leans against a wall at a 62 degree angle of elevation. Find the distance from the ladder's base to the wall

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\cos(62)=\frac{b}{12}\] \[b=12\cos(62)\] then a calculator

OpenStudy (anonymous):

cos 62 = adj/ hypo = b/12 => b = 12* cos62 = 12 * 1.88 = 22.6 ft